OpenStudy (anonymous):

Can some one help me please Evaluate the integral 25x^7 dx

hartnn (hartnn):

yes, firstly, 25 is a constant you can take it out of the integral, \(\huge =25\int x^7dx \) now, use the standard formula for integration of x^n dx \(\huge \int x^n dx=\dfrac{x^{n+1}}{n+1}+c\) here, put x =7, what u get ?

hartnn (hartnn):

**put n=7

OpenStudy (anonymous):

so would it be \[25*x^8/8\]

hartnn (hartnn):

yes, it will be, just add +c to it.

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

welcome ^_^